OpenStudy (anonymous):

(-2)2 - 4(-2)= the answer on this review came up as 12 not -12 can someone explain this one to me?

OpenStudy (mathteacher1729):

\[(-2)^2-4\cdot(-2)\] or \[(-2)\cdot2-4\cdot(-2)\]

OpenStudy (anonymous):

-2 times 2-4 times -2

OpenStudy (anonymous):

misprint (and where does -12 come into it?)

OpenStudy (anonymous):

wow lol yea -12 shouldve been an idea but why isnt it 4?

OpenStudy (tad1):

-4+8 = 4